# Optimizing Parameters for Static Equilibrium of Discrete Elastic Rods with Active-Set Cholesky

### v1

@@ VI-E Box-Constrained Quadratic Program Solver Comparisons
Figure 8: Log-scale profiles of residuals over time with various preconditioning techniques in MPRGP. The numbers in the last parentheses represent MPRGP iteration counts. Our ASC-MPRGP converges 18×18$\times$18 × faster than IC-MPRGP.
VI-E1 MPRGP Preconditioners
@@
    SAAMG-MPRGP: MPRGP with smoothed aggregation algebraic multigrid (SAAMG) preconditioning [46];
@@
Figure 8 shows log-scale profiles of convergence over time. Due to the stiffness of our system, MPRGP without preconditioning failed to converge within 10,000 iterations and was terminated. While WJ-MPRGP and SSOR-MPRGP effectively reduce the number of MPRGP iterations required, their preconditioning costs are non-negligible, and D-MPRGP performs slightly faster. Although SAAMG-MPRGP can be effective for systems with an M-matrix [46], it underperforms on our non-M-matrix system. By contrast, Cholesky-based preconditioners are particularly effective because they exploit the banded structure of the system with minimal overhead for a single factorization per MPRGP solve. IC-MPRGP requires significantly fewer iterations than D-MPRGP, and its moderate preconditioning cost results in much faster overall performance. Given the nearly fully filled band, MIC-MPRGP performs almost identically to IC-MPRGP. Our ASC-MPRGP, which employs full Cholesky factorization, benefits from the complete Cholesky factors without the limitations of (M)IC-MPRGP, which uses incomplete Cholesky with a GS fallback to avoid breakdown [44]. Consequently, while IC-MPRGP converges in 234 iterations taking 5.4ms, our ASC-MPRGP converges in just 19 iterations taking 0.3 ms, achieving around 18×18$\times$18 × faster performance.
Figure 9: Log-scale profiles of residuals over time with various BCQP solvers and permitted curvature change $\mu$. The numbers within the last parentheses indicate solver iteration counts. Our ASC-MPRGP outperforms IPM for $0.4\leq\mu$.
@@ VI-E2 BCQP Solvers
In this comparison, we exclude the penalty-based approach as it cannot accurately enforce the box constraints (comparisons between the penalty-based and active-set approaches are given in Sec. VI-C). Figure 9 shows log-scale profiles of convergence over time, and Table I summarizes the performance numbers (excluding PGS).
@@
Figure 10: Hair simulation with complex strand geometry. Hair strands significantly sag with naive initialization. Rest shape only optimization suffers from some sagging and unnatural hair lifting. Our method successfully achieves static equilibrium, exhibits natural motions due to the prescribed movements of the root vertices, and then restores the hairstyle to a form closely resembling the original.
VI-F Complex Strand Geometry
@@
Rest shape only optimization used 2.0 Newton iterations (with frequent failures in backtracking line search, leading to early termination) and 2.5 MPRGP iterations per solve on average (occasionally terminating at 100 iterations due to convergence issues), taking 0.6 s for the entire optimization process, whereas our method used 148.8 Newton iterations and 1.2 MPRGP iterations, taking 84.4 s. Although our method incurs higher costs than the rest shape only optimization (which fails to converge, thus suffering from sagging, unnatural lifting, and stability problems), we believe that the achieved static equilibrium justifies this extra one time initialization cost for forward simulations (which take 2.0 s per frame).
VII Conclusions and Future Work
@@
VII-A Undesirable Local Minima
@@
While our method ensures that strands achieve static equilibrium, certain perturbations may cause them to converge to other local minima, which may not align with user expectations. Although our solver enforces rest shape changes within specified box constraints to mitigate the risk of encountering such local minima, determining optimal values for these constraints can be challenging. The desirability of the resulting behaviors is often subjective, and the optimal values may vary depending on strand geometry and material stiffness. To accommodate diverse scenarios, it could be advantageous to assign different values of $\mu$ for each strand.
@@
In contrast to optimization methods focused solely on rest shape [5], our approach allows for modifications to material stiffness, which can be perceived as undesirable. While $\mathbf{W}$ in (9) assists in balancing changes between rest shape and material stiffness, a potentially more effective strategy may involve conducting separate and iterative parameter optimizations for rest shape and material stiffness, similar to block coordinate descent. Furthermore, to ensure smoothly varying material stiffness, it may be beneficial to employ a Laplacian-based regularizer or, alternatively, to optimize a limited set of representative stiffness parameters, which has the added advantage of reducing memory usage and optimization costs.
VII-B Active-Set Cholesky Preconditioner
@@
Our ASC preconditioner is applicable not only to banded systems but to other SPD ones too, and thus it is worth evaluating our preconditioner with stiff BCQPs. In particular, it should be able to support tree structures forming systems with block matrices (instead of banded systems) [50] without introducing any fill-in at off-diagonal blocks between the block matrices. It also seems promising to extend our preconditioner to (incomplete) LDLT factorization [54, 45] and to explore symmetric indefinite solvers that can handle box constraints. With small $\mu$, many box constraints can be activated, necessitating additional MPRGP iterations to manage active sets while failing to fully leverage our ASC preconditioning. In such cases, exploring active-set-free approaches, such as interior point methods [15, 52], may be promising. Although parallel execution over strands rendered extensive parallelization unnecessary in our framework, it would be valuable to optimize our preconditioner for parallel execution to fully utilize many-core architectures, particularly for strands discretized with a large number of vertices.
VII-C More General Inverse Problems
@@
Supporting anisotropic and inhomogeneous strands (e.g., with spatially varying density and radius), two-end clamping, and frictional contacts would be useful [30, 4]. To predict strand dynamics over time, developing differentiable physics approaches, such as those employing the adjoint method [2, 12], appears promising [55]. Extending our method to applications involving 2D/3D structured materials would also be of interest. While our method is guaranteed to reach static equilibrium at solver convergence, convergence conditions may depend on strand geometry and properties; thus, tuning solver parameters may be necessary in challenging scenarios. Additionally, solving the optimization problem without primal-dual decoupling could enhance the likelihood of convergence.

### v2

@@ VII-D Box-Constrained Quadratic Program Solver Comparisons
Figure 6: Log-scale profiles of residuals over time with various preconditioning techniques in MPRGP. The numbers in the last parentheses represent MPRGP iteration counts. Our ASC-MPRGP converges $18\times$ faster than IC-MPRGP.
VII-D1 MPRGP Preconditioners
@@
    SAAMG-MPRGP: MPRGP with SAAMG preconditioning [44];
@@
Figure 6 shows log-scale profiles of convergence over time. Due to the stiffness of our system, MPRGP without preconditioning failed to converge within 10,000 iterations and was terminated. While WJ-MPRGP and SSOR-MPRGP effectively reduce the number of MPRGP iterations required, their preconditioning costs are non-negligible, and D-MPRGP performs slightly faster. Although SAAMG-MPRGP can be effective for systems with an M-matrix [44], it underperforms on our non-M-matrix system. By contrast, Cholesky-based preconditioners are particularly effective because they exploit the banded structure of the system with minimal overhead for a single factorization per MPRGP solve. IC-MPRGP requires significantly fewer iterations than D-MPRGP, and its moderate preconditioning cost results in much faster overall performance. Given the nearly fully filled band, MIC-MPRGP performs almost identically to IC-MPRGP. Our ASC-MPRGP, which employs full Cholesky factorization, benefits from the complete Cholesky factors without the limitations of (M)IC-MPRGP, which uses incomplete Cholesky with a GS fallback to avoid breakdown [42]. Consequently, while IC-MPRGP converges in 234 iterations taking 5.4 ms, our ASC-MPRGP converges in just 19 iterations taking 0.3 ms, achieving around $18\times$ faster performance.
Figure 7: Log-scale profiles of residuals over time with various BCQP solvers and permitted curvature change $\mu$. The numbers in the last parentheses indicate solver iteration counts. Our ASC-MPRGP outperforms IPM for $\mu\geq 0.4$.
@@ VII-D2 BCQP Solvers
In this comparison, we exclude the penalty-based approach as it cannot accurately enforce the box constraints (comparisons between the penalty-based and active-set approaches are given in Sec. VII-B). Figure 7 shows log-scale profiles of convergence over time, and Table I summarizes the performance numbers (excluding PGS).
@@
Figure 8: Hair simulation with complex strand geometry. Hair strands significantly sag with naive initialization. Rest shape only optimization suffers from some sagging and unnatural hair lifting. Our method successfully achieves static equilibrium, exhibits natural motions driven by the prescribed movements of the root vertices, and then restores the hairstyle to a form closely resembling the original.
VII-E Complex Strand Geometry
@@
Rest shape only optimization used 2.0 Newton iterations (with frequent failures in backtracking line search, leading to early termination) and 2.5 MPRGP iterations per solve on average (occasionally terminating at 100 iterations due to convergence issues), taking 0.6 s for the entire optimization process, whereas our method used 148.8 Newton iterations and 1.2 MPRGP iterations, taking 84.4 s. Although our method incurs higher costs than the rest shape only optimization (which fails to converge, thus suffering from sagging, unnatural lifting, and stability problems), we believe that the achieved static equilibrium justifies this one time initialization cost for forward simulations (which take 2.0 s per frame).
VIII Conclusions and Future Work
@@
VIII-A Undesirable Local Minima
@@
While our method ensures that strands achieve static equilibrium, certain perturbations may cause them to fall into other local minima in forward simulations, which may not align with user expectations. Although our solver enforces rest shape changes within specified box constraints to mitigate the risk of encountering such local minima, determining optimal values for these constraints can be challenging. The desirability of the resulting behaviors is often subjective, and the optimal values may vary depending on strand geometry and material stiffness. To accommodate diverse scenarios, it could be advantageous to assign different values of $\mu$ for each strand.
@@
In contrast to optimization methods focused solely on the rest shape parameters [5], our approach allows for modifications to the material stiffness, which can be perceived as undesirable. While $\mathbf{W}$ in (5) assists in balancing changes between the rest shape and material stiffness parameters, a potentially more effective strategy may involve conducting separate and iterative parameter optimizations for the rest shape and material stiffness, similar to block coordinate descent. Furthermore, to ensure smoothly varying material stiffness, it may be beneficial to employ a Laplacian-based regularizer or, alternatively, to optimize a limited set of representative stiffness parameters, which has the added advantage of reducing memory usage and optimization costs.
VIII-B Active-Set Cholesky Preconditioner
@@
As our ASC preconditioner is based on a Cholesky-based direct solver which is particularly effective for stiff SPD linear systems, it is worth evaluating our preconditioner with stiff BCQPs. In particular, our preconditioner should work with tree structures which form block-structured systems (instead of banded ones) without introducing any fill-in at off-diagonal blocks between the block matrices [53]. It also seems promising to extend our preconditioner to (incomplete) LDLT factorization [57, 43] and to explore symmetric indefinite solvers that can handle box constraints. With small $\mu$, many box constraints can be activated, necessitating additional MPRGP iterations to manage active sets while failing to fully leverage our ASC preconditioning. In such cases, exploring active-set-free approaches, such as interior point methods [15, 55], may be promising. Although parallel execution over strands rendered extensive parallelization unnecessary in our framework, it would be valuable to optimize our preconditioner for parallel execution to fully utilize many-core architectures, particularly for strands discretized with a large number of vertices.
VIII-C More General Inverse Problems
@@
Supporting anisotropic and inhomogeneous strands, two-end clamping, and frictional contacts would be useful. In particular, incorporating contact forces in the parameter optimization would prevent extra stiffening and rest shape changes, giving more plausible results (e.g., for hairs on a head) [30, 3, 4]. To predict strand dynamics over time, developing differentiable physics approaches, such as those employing the adjoint method [2, 12], appears promising [58]. Extending our method to applications involving 2D/3D structured materials would also be of interest. While our method is guaranteed to reach static equilibrium at solver convergence, convergence conditions may depend on strand geometry and properties; thus, tuning solver parameters may be necessary in challenging scenarios. Additionally, solving the optimization problem without primal-dual decoupling could enhance the likelihood of convergence.
Acknowledgments
We thank the anonymous reviewers for their valuable suggestions, and Bo Yang, Qingyuan Zheng, Rundong Wu, and Yu Ju Chen for early discussions. We acknowledge the authors of [56] for publicly releasing the hair dataset. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2021-02524).